OpenStudy (anonymous):

someone please help me :( what is the quotient of the complex numbers below? 3+2i/1-5i A. -24 -10i B. -7/26 +17/26i C. 3-2/5i D. 13-13i

OpenStudy (anonymous):

multiply top and bottom be the conjugate of \(1-5i\)

OpenStudy (anonymous):

the conjugate of \(a+bi\) is \(a-bi\) and this works because \[(a+bi)(a-bi)=a^2+b^2\] a real number

OpenStudy (anonymous):

all the work is multiplyiing in the numerator \[\frac{3+2i}{1-5i}\times \frac{1+5i}{1+5i}\] \[=\frac{(3+2i)(1+5i)}{1^2+5^2}\]

OpenStudy (anonymous):

you know how to multiply complex numbers?

OpenStudy (anonymous):

no i don't, sadly.

OpenStudy (anonymous):

you have a choice

OpenStudy (anonymous):

you can just multiply out like you would say \((3+2x)(1+5x)\) you know how to do that?

OpenStudy (anonymous):

lol it is not that sad, you can learn it guess that is the point of this problem, but you can't divide if you don't know how to multiply for sure!

OpenStudy (anonymous):

can you multiply this\[(3+2x)(1+5x)\]

OpenStudy (anonymous):

you would combine like terms ? wouldn't you ?

OpenStudy (anonymous):

you ever had a math teacher that said "foil" or sommat?

OpenStudy (anonymous):

oh yes i just learned that, so 10x^2 + 17x + 3?

OpenStudy (anonymous):

ok yes so if you do it with \(i\) instead of \(x\) you get \[10i^2+17i+3\]but there is one difference

OpenStudy (anonymous):

since \(i=\sqrt{-1}\) you know \(i^2=-1\) so \[10i^2+17i+3=-10+17i+3=-7+17i\]

OpenStudy (anonymous):

and don't forget your denominator, which is \(1^2+5^2=1+25=26\)

OpenStudy (anonymous):

giving you \[\frac{-7+17i}{26}\] or \[\frac{-7}{26}+\frac{17}{26}i\]

OpenStudy (anonymous):

ok ok but i got so lost on that last step :(

OpenStudy (anonymous):

the \(-1\) part?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it is the definition of the symbol (letter) \(i\) used here it is the number whose square is \(-1\) i.e. by definition \(i^2=-1\)

OpenStudy (anonymous):

that is why \[10i^2+17i+3\] becomes \[-10+17i+3\]

OpenStudy (anonymous):

oh so only i^2 = -1, not i ?

OpenStudy (anonymous):

no the \(i\) stays there

OpenStudy (anonymous):

if \(i\) was \(-1\) no one would write \(i\) they would just write \(-1\) instead

OpenStudy (anonymous):

ok ok i get it now (: thank you !!! very helpful

OpenStudy (anonymous):

yw